OpenStudy (anonymous):

Give an example of two irrational numbers whose product is an irrational number

OpenStudy (anonymous):

@abbylei777

OpenStudy (anonymous):

2*sqrt2 is an irrational number Their product sqrt2*2*sqrt2 = 4 is a rational number

OpenStudy (anonymous):

I need an irrational number as the answer

Parth (parthkohli):

2 * root 2 is irrational, but but 2 is not irrational. A better example would be root 2 and root 3.

OpenStudy (anonymous):

\[\sqrt{2}*\sqrt{3} = \sqrt{6}\]

OpenStudy (anonymous):

pi and the square root of 44